# Leveraging Large Language Models to Detect Missed Peephole Optimizations

### v1

@@ Abstract.
By replacing small, suboptimal instruction sequences within programs with a more efficient equivalent, peephole optimization can not only directly optimize code size and performance, but also potentially enables further transformations in the subsequent optimization pipeline. Although peephole optimization is a critical class of compiler optimizations, discovering new and effective peephole optimizations is challenging as the instruction sets can be extremely complex and diverse. Previous methods either do not scale well or can only capture a limited subset of peephole optimizations.
In this work, we leverage Large Language Models (LLMs) to detect missed peephole optimizations. We propose Lampo, a novel automated framework that synergistically combines the creative but unreliable code optimization ability of LLMs with rigorous correctness verification performed by translation validation tools, integrated in a feedback-driven iterative process.
Through a comprehensive evaluation within LLVM ecosystems, we show that Lampo can successfully detect up to 17 out of 25 previously reported missed optimizations in LLVM on average, and that 22 out of 25 can potentially be found by Lampo with different LLMs. For comparison, the state-of-the-art superoptimizer for LLVM, Souper, identified 15 of them. Moreover, within seven months of development and intermittent experiments, Lampo found 26 missed peephole optimizations, 15 of which have been confirmed and 6 already fixed. These results demonstrate Lampo’s strong potential in continuously detecting missed peephole optimizations.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Compilers
@@ 1. Introduction
Despite its importance, the development of a comprehensive, effective peephole optimizer remains a significant challenge. Unlike other optimizations, such as dead code elimination or loop-invariant code motion, which may be implemented with a single, unified algorithm, peephole optimizations are a collection of pattern-matching and rewriting rules. The complexity and vastness of modern instruction sets cause the number of potential optimization patterns to multiply rapidly, making it exceedingly difficult to develop a complete optimizer. This is why optimizers in widely used compilers like LLVM are continuously evolving, with new patterns being generalized or added (Menendez and Nagarakatte, 2016; ins, nd).
@@
Superoptimization: Given a sequence of instructions, a superoptimizer aims to find the optimal instruction sequence through stochastic search or program synthesis (Sasnauskas et al., 2017; Schkufza et al., 2013; Joshi et al., 2002; Bansal and Aiken, 2006; Buchwald, 2015; Bansal and Aiken, 2008). The resulting transformation can then be generalized and implemented as a peephole optimization. However, these approaches are typically computationally expensive, which constrains the code size they can process and limits them to supporting only a subset of the instruction set. For example, Souper, a synthesized superoptimizer for LLVM IR, does not support memory, floating-point, or vector instructions.
@@
Lampo.  This paper explores the feasibility of using Large Language Models (LLMs) to discover new peephole optimizations. While LLMs are highly effective at creative exploration and reasoning, they are prone to generating incorrect or “hallucinated” answers. Conversely, formal verification can guarantee the correctness of an optimization but is often unsuited for the creative task of searching for new ones. We propose Lampo, a novel technique that resolves this tension by combining the creative search capabilities of LLMs with the rigorous guarantees of formal verification. The Lampo framework operates as an automated discovery engine for missed peephole optimizations. Its workflow is built around three key components: an extractor, an LLM-based optimizer, and a verifier. First, the extractor analyzes programs to identify candidate instruction sequences for optimization. These candidates are then passed to the LLM-based optimizer, which suggests a potentially more efficient equivalent. The verifier then rigorously checks the proposed optimization for refinement and performance improvement. If the optimization is verified, it is saved for further investigation. If the verification fails, the verifier’s output is used as targeted feedback to the LLM, guiding the LLM to correct its proposal in a new attempt. This closed-loop process allows us to systematically and automatically find new, verified peephole optimizations.
@@
We thoroughly evaluated Lampo’s ability to find new peephole optimizations. We first curated a benchmark suite of 25 missed peephole optimizations that were recently reported in the LLVM repository. Our results show that Lampo can identify up to 22 of them, while Souper, the state-of-the-art superoptimizer, only detected 15. Furthermore, we tested Lampo’s capability to discover unreported, missed optimizations. Within seven months of development and intermittent testing, Lampo found 26 missed peephole optimizations, with 15 confirmed and 6 already fixed in LLVM. Notably, Souper fails to identify 7 of these confirmed cases. These results clearly demonstrate the strong potential of our approach, which combines the creative search of LLMs with the rigorous correctness of formal verification, to continuously find new optimizations as the reasoning abilities of LLMs advance.
@@
  • We present Lampo, a novel framework that synergistically combines the creative search capabilities of LLMs with the rigorous guarantees of formal verification, providing an automated solution to the long-standing challenge of discovering new peephole optimizations.
@@
  • We conducted a comprehensive evaluation of Lampo within the LLVM ecosystem, demonstrating its ability to find unreported missed optimizations. Lampo has successfully identified 26 missed peephole optimizations, with 15 confirmed and 6 already fixed in LLVM. Notably, 7 of these optimizations cannot be detected by Souper, the state-of-the-art superoptimizers for LLVM. These results highlight the advantages of Lampo over prior approaches and establish it as a validated, practical tool for compiler development.
@@ 2. Background
We built Lampo on top of LLVM and used Lampo to find missed peephole optimizations within LLVM. We chose LLVM because of its flexible and user-friendly tool-chain. While our implementation targets LLVM, the core concept of Lampo is general and can be applied to other compilers. This section introduces necessary background knowledge on LLVM IR, peephole optimizations, and translation validation.
@@ 2.1. LLVM Intermediate Representation
LLVM is an open-source compiler framework for various programming languages and architectures (Lattner, 2002; Lattner and Adve, 2004; llv, ndc). It has a low-level, typed IR in static single-assignment form, designed for expressiveness and extensibility.
LLVM IR programs consist of modules. Each module is a translation unit of the input program, and also is the top level container of all other IR objects, such as functions and global variables (llv, ndb). Figure 1(d) illustrates an LLVM IR module (irrelevant instructions, attributes and declarations are hidden for clarity) that contains a function generated from the original Rust function clamp in Figure 1(a). The IR function is organized into basic blocks, such as vector.body (line 4) shown in the figure. Each basic block comprises a sequence of instructions, has a single entry point (optionally denoted by a label), and ends with a terminator instruction that determines which basic block will be executed next, or whether the function will return. (e.g., br and ret instructions) (llv, ndb).
@@ 2.2. Peephole Optimization in LLVM
However, the development of a comprehensive and effective peephole optimizer is extremely challenging (llv, nda; Davidson and Fraser, 1984). Currently, the LLVM compiler contains over 25,000 lines of C++ code dedicated to peephole optimizations, and this number is still growing with new patterns added or existing patterns being generalized (Menendez and Nagarakatte, 2016; ins, nd).
@@
Figure 2. The overall workflow of Lampo. The verification phase is outlined by the dashed line.
@@ 2.4. Translation Validation
Alive2.Alive2 (Lopes et al., 2021) is an automated translation validation tool for LLVM. Given a pair of LLVM functions, Alive2 verifies whether the transformation from the source function to the target function is correct (i.e., whether the source is refined by the target). For example, given the LLVM function pair shown in Figures 1(b) and 1(c), Alive2 can prove that the transformation from src to tgt is correct. Conversely, if a transformation is proven to be incorrect, Alive2 provides a counterexample that refutes the refinement relation between the source and the target. Alive2 is widely used in the LLVM community and is integrated into the LLVM development workflow (llv, nda).
@@ 3. Methodology
Figure 2 shows the overall workflow of Lampo. Given a corpus of programs in LLVM IR that have been compiled and optimized by LLVM, Lampo first traverses the basic blocks in these IR programs and extracts all the dependent instruction sequences within each basic block (step 1). For each instruction sequence, Lampo prompts an LLM to output the optimal instruction sequence if the original instruction sequence is considered suboptimal (step 2). However, LLMs are not always reliable. The candidate produced by LLMs can, for example, (1) have syntax error, (2) be sub-optimal, (3) be not canonical, (4) have no refinement relation to the original instruction sequence, or (5) be not interesting as it is identical to the original sequence or is less efficient. To tackle this limitation of LLMs, we incorporate a verification phase as outlined by the dashed line to check all these mistakes that can be made by the LLM (steps 3, 4, and 5). If the instruction sequence proposed by the LLM has syntax error or the optimization is proven to be incorrect, Lampo uses the syntax error message or a counterexample provided by the correctness verifier as feedback, and prompts the LLM to output a correct version based on the feedback until the reattempt limit is reached (steps 6). In our implementation, we set the limit to two. Eventually, if the result generated by the LLMs is interesting (discussed in section 3.3) and the optimization is proven to be correct, Lampo saves the pair of instruction sequences as a potential missed optimization, which is then further analyzed and reported to developers (step 7). The remainder of this section details the workflow.
@@
(a) An instruction sequence (wrapped as a function) extracted by Lampo from the LLVM IR module shown in Figure 1(d). The dashed box highlights the instructions to be optimized.
@@
Figure 3. An illustrative example demonstrating the workflow of Lampo.
3.1. Illustrative Example
We use the real-world LLVM module in Figure 1(d) (section 2) to illustrate the workflow of Lampo. Given this module, Lampo first traverses all the basic blocks in each function of this module and extracts instruction sequences from them. Figure 3(a) shows one of the instruction sequences extracted from the basic block vector.body.
Lampo then provides the extracted instruction sequence to an LLM and prompts it to produce a correct and optimal instruction sequence. In this example, Lampo uses gemini-2.0-flash-thinking-exp-01-21 (referred to as Gemini2.0T) as its underlying LLM, and the initial candidate produced by this model, shown in Figure 3(b), contains a syntax error.
After obtaining this candidate, Lampo initiates the verification process. It first invokes the LLVM optimizer opt to check candidate’s syntax and to potentially further optimize and canonicalize the candidate. In this case, since the candidate contains a syntax error, opt outputs an error message as shown in Figure 3(c). Lampo then feeds this error message back to the Gemini2.0T, prompting it to produce a new candidate. Figure 3(d) shows the corrected candidate generated by Gemini2.0T after receiving the feedback.
With this new candidate, Lampo again invokes opt, and now the syntax check passes. Meanwhile, no further optimization and canonicalization is performed by opt since the candidate is already optimal and canonical. The subsequent interestingness checking (discussed in section 3.3) also passes, as the candidate contains one fewer instructions than the original instruction sequence. Finally, Alive2 formally confirms the correctness of the proposed optimization, and Lampo records the original instruction sequence and the candidate as a potential missed optimization.
@@ 3.2. Extracting Instruction Sequences
Motivation.  The motivation for extracting instruction sequences from an entire LLVM IR module is threefold. First, an LLVM IR module can be extremely large and complex, and from our experience, existing LLMs struggle to identify missed peephole optimizations directly from a complete module. By prompting LLMs to focus on optimizing small instruction sequences, we increase the likelihood of discovering potential missed optimizations. Second, extracting instruction sequences from each IR module allows Lampo to exclude irrelevant information and deduplicate sequences that appear multiple times in the corpus, and thereby increase the overall efficiency and save the cost for LLM inference. Finally, slicing modules into instruction sequences facilitates subsequent verification and analysis by restricting the scope to a small instruction sequence in each iteration.
@@
6     foreach seq $\in$ seq_set do
7        wrapped_func $\leftarrow$ WrapAsFunc(seq)

### v2

# LPO: Discovering Missed Peephole Optimizations with Large Language Models

@@ Abstract.
Peephole optimization is an essential class of compiler optimizations that targets small, inefficient instruction sequences within programs. By replacing such suboptimal instructions with refined and more optimal sequences, these optimizations not only directly optimize code size and performance, but also enable more transformations in the subsequent optimization pipeline. Despite their importance, discovering new and effective peephole optimizations remains challenging due to the complexity and breadth of instruction sets. Prior approaches either lack scalability or have significant restrictions on the peephole optimizations that they can find.
This paper introduces LPO, a novel automated framework to discover missed peephole optimizations. Our key insight is that, Large Language Models (LLMs) are effective at creative exploration but susceptible to hallucinations; conversely, formal verification techniques provide rigorous guarantees but struggle with creative discovery. By synergistically combining the strengths of LLMs and formal verifiers in a closed-loop feedback mechanism, LPO can effectively discover verified peephole optimizations that were previously missed.
We comprehensively evaluated LPO within LLVM ecosystems. Our evaluation shows that LPO can successfully identify up to 22 out of 25 previously reported missed optimizations in LLVM. In contrast, the recently proposed superoptimizers for LLVM, Souper and Minotaur detected 15 and 3 of them, respectively. More importantly, within eleven months of development and intermittent testing, LPO found 62 missed peephole optimizations, of which 28 were confirmed and an additional 13 had already been fixed in LLVM. These results demonstrate LPO’s strong potential to continuously uncover new optimizations as LLMs’ reasoning improves.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790184isbn: 979-8-4007-2359-9/2026/03ccs: Software and its engineering Compilers
@@ 1. Introduction
Despite its importance, the development of a comprehensive, effective peephole optimizer remains a significant challenge. Unlike other optimizations, such as dead code elimination or loop-invariant code motion, which may be implemented with a single, unified algorithm, peephole optimizations are a collection of pattern-matching and rewriting rules. The complexity and vastness of modern instruction sets cause the number of potential optimization patterns to multiply rapidly, making it exceedingly difficult to develop a complete optimizer. This is why optimizers in widely used compilers like LLVM are continuously evolving, with new patterns being generalized or added (D. Menendez and S. Nagarakatte (2016); 19).
@@
Superoptimization: Given a sequence of instructions, a superoptimizer aims to find the optimal instruction sequence through stochastic search or program synthesis (Sasnauskas et al., 2017; Schkufza et al., 2013; Joshi et al., 2002; Bansal and Aiken, 2006; Buchwald, 2015; Bansal and Aiken, 2008; Liu et al., 2024). The resulting transformation can then be generalized and implemented as a peephole optimization. However, these approaches are typically computationally expensive, which constrains the code size they can process and limits them to supporting only a subset of the instruction set. For example, Souper, a synthesized superoptimizer for LLVM IR, does not support memory, floating-point, or vector instructions.
@@
LPO.  This paper explores the feasibility of using Large Language Models (LLMs) to discover new peephole optimizations. While LLMs are highly effective at creative exploration and reasoning, they are prone to generating incorrect or “hallucinated” answers. Conversely, formal verification can guarantee the correctness of an optimization but is often unsuited for the creative task of searching for new ones. We propose LPO, a novel technique that resolves this tension by combining the creative search capabilities of LLMs with the rigorous guarantees of formal verification. The LPO framework operates as an automated discovery engine for missed peephole optimizations. Its workflow is built around three key components: an extractor, an LLM-based optimizer, and a verifier. First, the extractor analyzes programs to identify candidate instruction sequences for optimization. These candidates are then passed to the LLM-based optimizer, which suggests a potentially more efficient equivalent. The verifier then rigorously checks the proposed optimization for refinement and performance improvement. If the optimization is verified, it is saved for further investigation. If the verification fails, the verifier’s output is used as targeted feedback to the LLM, guiding the LLM to correct its proposal in a new attempt. This closed-loop process allows us to systematically and automatically find new, verified peephole optimizations.
@@
We thoroughly evaluated LPO’s ability to find new peephole optimizations. We first curated a benchmark suite of 25 missed peephole optimizations that were recently reported in the LLVM repository. Our results show that LPO can identify up to 22 of them, while Souper and Minotaur, two recently proposed synthesizing superoptimizers, detected 15 and 3. Furthermore, we tested LPO’s capability to discover unreported, missed optimizations. Within eleven months of development and intermittent testing, LPO found 62 missed peephole optimizations, with 28 confirmed and an additional 13 already fixed in LLVM. Notably, Souper and Minotaur fail to identify 26 and 31 of these confirmed or fixed cases, respectively. These results clearly demonstrate the strong potential of our approach, which combines the creative search of LLMs with the rigorous correctness of formal verification, to continuously find new optimizations as the reasoning abilities of LLMs advance.
@@
  • We present LPO, a novel framework that synergistically combines the creative search capabilities of LLMs with the rigorous guarantees of formal verification, providing an automated solution to the long-standing challenge of discovering new peephole optimizations.
@@
  • We comprehensively evaluated LPO in the LLVM ecosystem and showed that it can uncover previously unreported missed optimizations. LPO identified 62 missed peephole optimizations, with 28 confirmed and 13 already fixed in LLVM. Notably, 26 and 31 of these optimizations cannot be detected by Souper and Minotaur, two synthesizing superoptimizers for LLVM, respectively. These results highlight the advantages of LPO over prior approaches and establish it as a validated, practical tool for compiler development.
  • To ensure reproducibility and facilitate future research, we have released the replication package and experimental data at https://github.com/uw-pluverse/lpo-artifact.
@@ 2. Background
We built LPO on top of LLVM and used LPO to find missed peephole optimizations within LLVM. We chose LLVM because of its flexible and user-friendly tool-chain. While our implementation targets LLVM, the core concept of LPO is general and can be applied to other compilers. This section introduces necessary background knowledge on LLVM IR, peephole optimizations, and translation validation.
@@ 2.1. LLVM Intermediate Representation
LLVM is an open-source compiler framework for various programming languages and architectures (C. A. Lattner (2002); C. Lattner and V. Adve (2004); 35). It has a low-level, typed IR in static single-assignment form, designed for expressiveness and extensibility.
LLVM IR programs consist of modules. Each module is a translation unit of the input program, and also is the top level container of all other IR objects, such as functions and global variables (20). Figure 1(d) illustrates an LLVM IR module (irrelevant instructions, attributes and declarations are hidden for clarity) that contains a function generated from the original Rust function clamp in Figure 1(a). The IR function is organized into basic blocks, such as vector.body (line 4) shown in the figure. Each basic block comprises a sequence of instructions, has a single entry point (optionally denoted by a label), and ends with a terminator instruction that determines which basic block will be executed next, or whether the function will return. (e.g., br and ret instructions) (20).
@@ 2.2. Peephole Optimization in LLVM
However, the development of a comprehensive and effective peephole optimizer is extremely challenging (12; J. W. Davidson and C. W. Fraser (1984)). Currently, the LLVM compiler contains over 25,000 lines of C++ code dedicated to peephole optimizations, and this number is still growing with new patterns added or existing patterns being generalized (D. Menendez and S. Nagarakatte (2016); 19).
@@ 2.4. Translation Validation
Alive2.Alive2 (Lopes et al., 2021) is an automated translation validation tool for LLVM. Given a pair of LLVM functions, Alive2 verifies whether the transformation from the source function to the target function is correct (i.e., whether the source is refined by the target). For example, given the LLVM function pair shown in Figures 1(b) and 1(c), Alive2 can prove that the transformation from src to tgt is correct. Conversely, if a transformation is proven to be incorrect, Alive2 provides a counterexample that refutes the refinement relation between the source and the target. Alive2 is widely used in the LLVM community and is integrated into the LLVM development workflow (12).
@@ 3. Methodology
Figure 2. The overall workflow of LPO. The verification phase is outlined by the dashed line.
Figure 2 and Algorithm 1 show the overall workflow of LPO.
Step 1: Extract instruction sequences.  Given a corpus of programs in LLVM IR that have been compiled and optimized by LLVM, LPO first traverses the basic blocks in these IR programs and extracts all the dependent instruction sequences within each basic block (Algorithms 11). The details about the extraction process is introduced in section 3.2.
Step 2 3 7: Optimize the instruction sequence.  For each instruction sequence, LPO prompts an LLM to output the optimal instruction sequence (Algorithm 1 and step 2). For the candidate generate by the LLM, LPO runs three checks to verify whether it potentially reveals a valid optimization (Algorithms 11). If all checks are passed, the instruction sequence and the corresponding candidate generated by the LLM are saved for further analysis (Algorithm 1 and step 7).
Step 3 4 5 6: Verify the candidate and provide feedback.  LLMs are not always reliable. The candidate produced by LLMs can, for example, (1) have syntax error, (2) be sub-optimal, (3) be not canonical, (4) have no refinement relation to the original instruction sequence, or (5) be not interesting as it is identical to the original sequence or is less efficient. To tackle this limitation of LLMs, LPO incorporates a verification phase as outlined by the dashed line in Figure 2 to check all these mistakes that can be made by the LLM (steps 3, 4, and 5). First, the instruction sequence is sent to the LLVM optimizer, opt. If there is any syntax error, LPO uses the error message from opt as feedback and prompt the LLM to generate a new candidate (Algorithms 11). Otherwise, the optimized instruction sequence output by opt becomes the new candidate (Algorithm 1). Next, LPO checks whether the candidate is interesting, i.e., whether the candidate potentially manifests a beneficial optimization (more details in section 3.3). If the candidate is not interesting (e.g., is identical to the original instruction sequence), LPO abandons this instruction sequence and moves to the next one (Algorithms 11). If the candidate is interesting, LPO further checks the correctness with Alive2 (Algorithm 1). If the optimization is proven to be incorrect, LPO uses the counterexample output by Alive2 as the feedback and prompts the LLM to output a correct version based on the feedback (Algorithms 11). A parameter ATTEMPT_LIMIT can be configured to specify how many attempts are allowed at most for optimizing each instruction sequence (Algorithm 1). In our implementation, this parameter is set to 2.
The remainder of this section details the workflow. Figure 3 illustrates the workflow with a concrete example. section 3.2 elaborates on how LPO extracts instruction sequences from the corpus. section 3.3 details the verification process.
Input : corpus, a list of LLVM IR module
Output : A set of instruction sequence pairs that potentially reveal missed peephole optimizations
1 instSeqs $\leftarrow\varnothing$
2 dedup_set $\leftarrow\varnothing$
3 foreach module $\in$ corpus do
4  instSeqs $\leftarrow$ instSeqs $\cup$ Extract(module, dedup_set)
5
6 foreach instSeq $\in$ instSeqs do
7  counter $\leftarrow$ 0
8  feedback $\leftarrow$ empty string
9  while counter ¡ ATTEMPT_LIMIT do
10     candidate $\leftarrow$ InvokeLLM(instSeq, feedback)
11     result_OPT $\leftarrow$ RunOPT(candidate)
12     if result_OPT.is_failed then
13        counter $\leftarrow$ counter + 1
14        feedback $\leftarrow$ result_OPT.error_message
15        continue
16
17     candidate $\leftarrow$ result_OPT.new_candidate
[…]
19
20     result_Alive2 $\leftarrow$ RunAlive2(instSeq, candidate)
[…]
22        counter $\leftarrow$ counter + 1
23        feedback $\leftarrow$ result_Alive2.counter_example
24        continue
25
26     save instSeq and candidate for further analysis
27     break
28
29
Algorithm 1 The overall workflow of LPO
3.1. Illustrative Example
@@
(a) An instruction sequence (wrapped as a function) extracted by LPO from the LLVM IR module shown in Figure 1(d). The dashed box highlights the instructions to be optimized.
@@
Figure 3. An illustrative example demonstrating the workflow of LPO.
We use the real-world LLVM module in Figure 1(d) (section 2) to illustrate the workflow of LPO. Given this module, LPO first traverses all the basic blocks in each function of this module and extracts instruction sequences from them. Figure 3(a) shows one of the instruction sequences extracted from the basic block vector.body.
LPOthen provides the extracted instruction sequence to an LLM and prompts it to produce a correct and optimal instruction sequence. In this example, LPO uses gemini-2.0-flash-thinking-exp-01-21 (referred to as Gemini2.0T) as its underlying LLM, and the initial candidate produced by this model, shown in Figure 3(b), contains a syntax error.
After obtaining this candidate, LPO initiates the verification process. It first invokes the LLVM optimizer opt to check candidate’s syntax and to potentially further optimize and canonicalize the candidate. In this case, since the candidate contains a syntax error, opt outputs an error message as shown in Figure 3(c). LPO then feeds this error message back to the Gemini2.0T, prompting it to produce a new candidate. Figure 3(d) shows the corrected candidate generated by Gemini2.0T after receiving the feedback.
With this new candidate, LPO again invokes opt, and now the syntax check passes. Meanwhile, no further optimization and canonicalization is performed by opt since the candidate is already optimal and canonical. The subsequent interestingness checking (discussed in section 3.3) also passes, as the candidate contains one fewer instructions than the original instruction sequence. Finally, Alive2 formally confirms the correctness of the proposed optimization, and LPO records the original instruction sequence and the candidate as a potential missed optimization.
@@ 3.2. Extracting Instruction Sequences
Motivation.  The motivation for extracting instruction sequences from an entire LLVM IR module is threefold. First, an LLVM IR module can be extremely large and complex, and from our experience, existing LLMs struggle to identify missed peephole optimizations directly from a complete module. By prompting LLMs to focus on optimizing small instruction sequences, we increase the likelihood of discovering potential missed optimizations. Second, extracting instruction sequences from each IR module allows LPO to exclude irrelevant information and deduplicate sequences that appear multiple times in the corpus, and thereby increase the overall efficiency and save the cost for LLM inference. Finally, slicing modules into instruction sequences facilitates subsequent verification and analysis by restricting the scope to a small instruction sequence in each iteration.
@@
6     foreach instSeq $\in$ seq_set do
7        wrapped_func $\leftarrow$ WrapAsFunc(instSeq)